1.Assume you need to evaluate an ecotouristic project. Each year the project requires maintenance costs equal to $200 and you h2.Assume you have certified coffee plantation project. Each year you would have variable costs equal to $250 and benefits of $5
Answer #1

Q1.

Given,

(Note that costs are negative flows and benefits are positive flows)

Net cash flow in year 1 = -$1,000

Net cash flows in years 2 through 5 = -$200 + $500 = $300

Market Interest rate = 5% per year

From the compound interest table, we obtain

(P/F, 5%, 1) = 0.9524

(P/A, 5%, 4) = 3.546

Net Present Value of the above cash flow = Net Cash flow in year 1 x (P/F, 5%, 1) + Net cash flow in year 2 through 5 x (P/A, 5%, 4) x (P/F, 5%, 1) = -$1,000 x 0.9524 + $300 x 3.546 x 0.9524 = -$952.4 + $1,013.2 = $60.8

Ans: $60.8

Q2.

Net cash flow in year 1 = -$800

Net cash flows in years 2 through 5 = -$250 + $500 = $250

Net Present Value of the above cash flow = Net Cash flow in year 1 x (P/F, 5%, 1) + Net cash flow in year 2 through 5 x (P/A, 5%, 4) x (P/F, 5%, 1) =-$800 x 0.9524 + $250 x 3.546 x 0.9524 = -$761.92 + $844.30 = $82.38

Ans: $82.38